OpenStudy (anonymous):

Describe each of the following set by listing the elements in the set: {x ∈ ℚ | x(x2 – 2)(2x + 3) = 0} {-15, -5, -3, -1, 1, 3, 5, 15} {-2, -3/2, 0, 2} {} {-3/2, 0}

geerky42 (geerky42):

Just start by solving \(x(x^2-2)(2x-3)=0\)

geerky42 (geerky42):

If \(ab = 0\), then either \(a=0\) or \(b=0\). You can apply same logic here

OpenStudy (anonymous):

how come people are one on the weekend?? everyone needs a mega break from learning,i just logged on tho check my page. seams stress full.

OpenStudy (dayakar):

x=0 or \[x^{2}-2=0\] \[x^{2}=2\]\[x = \sqrt{2}or-\sqrt{2}\]

OpenStudy (dayakar):

therefore \[x={0.\sqrt{2},-\sqrt{2},3/2}\]

OpenStudy (dayakar):

verify ur options

geerky42 (geerky42):

I mean \(x(x^2-2)(2x+3)=0\)

OpenStudy (anonymous):

@dayakar I can see why you may have started with x=0; but, why did you also do x^2-2=0?

geerky42 (geerky42):

So solutions would be \(0, ~\text-\sqrt2, \sqrt2, \text-3/2\) Now check which elements are in set \(\mathbb Q\).

geerky42 (geerky42):

If \(x(\color{red}{x^2-2})(\color{blue}{2x+3})=0\), then \(x=0, ~\color{red}{x^2-2}=0, \text{ or }\color{blue}{2x+3}=0\)

OpenStudy (anonymous):

Okay, that makes sense. I get it now.

OpenStudy (anonymous):

and sqrt can't be expressed as a quotient. so it's just 0,-3/2

geerky42 (geerky42):

You are correct.

OpenStudy (anonymous):

Thank you for the help. I just became your fan :)

geerky42 (geerky42):

No problem :)